how do you solve these 2 problems.
9-6^2/4+12*4/2^4-3
789.3+4.287+13+2.9

Answer by carter6949(26)

to simplify all you do is solve the equation
here something to help you remember the steps
Please - P = parentheses
Excuse - E = exponents
My - M = multipication
Dear - D = division
Aunt - A = addition
Sally - S = subtraction
---------------------------------------------------------
Now lets look at the first equation
9-6^2/4+12*4/2^4-3
9-36/4+12*4/16-3 first you will solve the parentheses (6^2=36)(2^4=16)
9-36/4+48/16-3 second you will solve the multiplication (12*4=18)
9-9+3-6 third you will divide (36/4=9)(48/16=3)
9-12-3 fourth you will add (9+3=12)
-3-3 and last you will subtract (-3-3=-6)
-->-6<--YOUR ANSWER
---------------------------------------------------------
Now lets look at the second equation
ok look at the PEMDAS do i have any parentheses? NO, do i have any exponents? NO, do i have any Multiplicaton or divison? NO, do i have any addition? YES
---
789.3+4.287+13+2.9 add these numbers together to get this answer
-->809.487<-- Your answer
Thank You for using Algebra.com. I am Tutor Thomas if you need any 1 on 1 live help online please feel free to email me at math1075@yahoo.com. I tutor for free using yahoo messenger. doodle. Thanks Again, Thomas